# Bruno Dewer

### v2

@@ 1. $\bullet$
    However, these two conditions also make sense for a normal, possibly singular, surface. As already mentioned in the introduction above, the surface $S$ we obtain as a general anticanonical divisor of a Gorenstein weighted projective $3$-space $\mathbf{P}$ is expected to have canonical singularities (to me more previse, ADE singularities). However, it is normal and satisfies the conditions $K_{S}=0$ and $h^{1}(S,\mathcal{O}_{S})=0$. As other authors did before, we refer to such a surface $S$ as a K3 surface with ADE singularities, and we sometimes simply say K3 surface.
@@ 2 The Gorenstein weighted projective spaces of dimension $3$
Remark. If $\mathbf{P}$ is not Gorenstein, then the canonical divisor $K_{\mathbf{P}}$ is not Cartier around at least one coordinate point, i.e., a point given by the vanishing of $3$ out of the $4$ homogeneous coordinates. As an example, consider $\mathbf{P}=\mathbf{P}(1,2,4,5)_{[u:v:s:t]}$, with $\mathcal{O}_{\mathbf{P}}(-K_{\mathbf{P}})=\mathcal{O}_{\mathbf{P}}(10)$. Since $4$ is not a divisor of $10$, then any local section of $\mathcal{O}_{\mathbf{P}}(10)$ around $p_{s}=[0:0:1:0]$ is a quotient of polynomials with numerator a sum of monomials, all of which admit at least one of the coordinates $u,v,t$ as a factor. Hence such a local section vanishes at $p_{s}$.
@@
Moreover, up to isomorphism, we may assume that any three of the four weights $a_{0},a_{1},a_{2},a_{3}$ are coprime. This condition is called well formedness.
@@ 2.2 The very ampleness of $-K_{\mathbf{P}}$
Let $S$ be a general member of $|-K_{\mathbf{P}}|$ and $\Gamma$ a general member of $|-K_{\mathbf{P}}|_{S}|$. We first prove that $-K_{\mathbf{P}}|_{S}$ is very ample. Thanks to Lemma 2.7, it is enough to prove that $H^{0}(S,\mathcal{O}_{S}(-K_{\mathbf{P}}|_{S}))^{\otimes n}\to H^{0}(S,\mathcal{O}_{S}(-nK_{\mathbf{P}}|_{S}))$ is onto for all $n\geq 1$. It is naturally the case for $n=1$. Assume now that $n>1$ and that it holds up to rank $n-1$. Consider the following commutative diagram.
@@
Moreover, we know that K3 surfaces with ADE singularities are projectively normal. Indeed, if $S$ is a singular K3 surface in $\mathbf{P}^{g}$, we may consider a resolution $S^{\prime}\to S\subset\mathbf{P}^{g}$ with $S^{\prime}$ a smooth K3 and apply [SD74, Theorem 6.1.$(ii)$]. It follows that the restriction map $H^{0}(\mathbf{P},\mathcal{O}_{\mathbf{P}}(-nK_{\mathbf{P}}))\to H^{0}(S,\mathcal{O}_{S}(-nK_{\mathbf{P}}|_{S}))$ is onto for all $n\geq 1$. Consider the following commutative diagram with $n>1$, whose right column is exact.
@@ 3 Birational models
The first step consists in the introduction of a suitable Veronese map on each $\mathbf{P}$, which is an embedding $v_{n}:\mathbf{P}\hookrightarrow X$ where $X$ is a weighted projective space of dimension $4$ so that the anticanonical model $\mathbf{P}\to\mathbf{P}^{g+1}$ factors as a composition
${\mathbf{P}}$${X}$${\mathbf{P}^{g+1}}$$\scriptstyle{v_{n}}$
where $X\dashrightarrow\mathbf{P}^{g+1}$ is a rational map that we will specify.
@@ 3.1 The Veronese maps
By a divisibility criterion, we can check fairly easily that $|\mathcal{O}_{X}(\frac{\sigma}{n})|$ is not always basepoint-free. This criterion is purely numerical: $|\mathcal{O}_{X}(\frac{\sigma}{n})|$ is basepoint-free if and only if $\frac{\sigma}{n}$ is divisible by all the weights of $X$. Namely, it is not the case for #10, #12 and #14, for which the induced map $X\dashrightarrow\mathbf{P}^{g+1}$ is nonregular.
@@ 3.2 The birational models
In all cases but #11, the image of $X$ in $\mathbf{P}^{g+1}$ is a cone with vertex a point which belongs to the image of $\mathbf{P}\to\mathbf{P}^{g+1}$. This follows from the fact that $\frac{\sigma}{n}$ equals the largest weight of $X$; say, $X=\mathbf{P}(d_{0},d_{1},d_{2},d_{3},d_{4})$ with $\frac{\sigma}{n}=d_{4}$. Then the map given by the linear system $|\mathcal{O}_{X}(\frac{\sigma}{n})|$ is the following.
$[x_{0}:x_{1}:x_{2}:x_{3}:x_{4}]\in X\mapsto[\mathtt{f}_{0}:\mathtt{f}_{1}:\cdots:\mathtt{f}_{g}:x_{4}]$
@@
$\varphi:[x:y:z:w]\in\mathbf{P}\mapsto[u_{0}:u_{1}:u_{2}:v]=[x^{5}:xy:z:w].$
@@
$u_{0}f_{4}(\mathbf{u},v)+u_{1}^{5}=g_{4}(\mathbf{u},v)=0$
with $g_{4}$ a general quartic on $\mathbf{P}^{\prime}$. This is summed up in the following commutative diagram.
@@ 4 Extensions of $\mathbf{P}$
Let $\Gamma\subset\mathbf{P}^{g-1}$ be a smooth canonical curve of genus $g\geq 11$ such that $\mathrm{Cliff}(\Gamma)\geq 3$. Let $\alpha=\mathrm{coker}(\Phi_{K_{\Gamma}})$ and $Y\subset\mathbf{P}^{g-1+\alpha}$ be a maximal extension of $\Gamma$ with $\dim Y\geq 2$. Then $Y$ is the universal extension of $\Gamma$ if and only if it contains no cone over $\Gamma$ as a linear surface section.
@@
Let us denote by $\mathcal{H}$ the family of $g$-planes $\Lambda\subset\mathbf{P}^{g-1+\alpha}$ such that $\Gamma\subset\Lambda$, and $\mathcal{S}=\mathbf{P}(\mathrm{coker}(\Phi_{K_{\Gamma}}))$ the space which parametrizes the surface extensions of $\Gamma$. Then we have $\mathcal{H}\simeq\mathcal{S}\simeq\mathbf{P}^{\alpha-1}$, and the map $\mathcal{H}\to\mathcal{S}$ which maps $\Lambda$ to $Y\cap\Lambda$ is well-defined since no linear surface section of $Y$ is a cone. Moreover, it lifts to a linear map $\mathbf{C}^{\alpha}\to\mathbf{C}^{\alpha}$ which has trivial kernel, i.e., it is an isomorphism (see [CD24, 6.6]).
It follows that $Y$ contains all the surface extensions of $\Gamma$ as unique linear sections. By Theorem 2.3, it is the universal extension of $\Gamma$. ∎
@@ 4.1 $\mathbf{P}=\mathbf{P}(1,4,5,10)$
The fivefold $Y$ is a maximal extension of $\Gamma$ by Theorem 2.3 since it has dimension $1+\alpha(\Gamma,K_{\Gamma})$ and it contains $\mathbf{P}=\mathbf{P}(1,4,5,10)$ as a $3$-fold linear section. Indeed, as indicated in Table 3, $\mathbf{P}$ embeds into $\mathbf{P}(1,1,1,2,4)_{[\mathbf{u}:v:s_{0}]}$ as the quintic hypersurface $u_{0}s_{0}=u_{1}^{5}$. It follows that $Y$ is an extension of $\mathbf{P}$, as
@@ 2. $(ii)$
which are all linear sections of $Y$ and contain $\Gamma$ as a hyperplane section. Indeed, the curve $\Gamma$ is cut out on all of them by $\left\{g_{4}(\mathbf{u},v)=0\right\}$. Among them, those that are members of the linear system $\mathcal{L}$ are those parameterized by $\lambda_{1}=\lambda_{2}=0$. The other surface extensions of $\Gamma$ could not be recovered from $\mathbf{P}$ and required the introduction of the birational model $\varphi:\mathbf{P}\dashrightarrow\mathbf{P}^{\prime}$.
@@ 4.2 $\mathbf{P}=\mathbf{P}(1,2,6,9)$
Besides, $Y$ is not a cone, by the same arguments as those mentioned in the proof of Lemma 4.2. The fourfold $Y$ is thus a maximal extension of $\Gamma$. Recall from Table 3 that $\mathbf{P}$ is the hypersurface of $\mathbf{P}(1,1,3,5,9)_{[\mathbf{u}:v:s:t_{0}]}$ given by the equation $u_{0}t_{0}=s^{2}$. This shows that $\mathbf{P}=Y\cap\left\{t_{1}=0\right\}$. ∎
@@ 4.3 $\mathbf{P}=\mathbf{P}(1,3,8,12)$
Hence it is a maximal extension of $\Gamma$. It is also an extension of $\mathbf{P}$; indeed, we know from Table 3 that $\mathbf{P}$ is the hypersurface $u_{0}t_{0}=v^{3}$ is $\mathbf{P}(1,1,3,4,8)_{[\mathbf{u}:v:s:t_{0}]}$. This exhibits $\mathbf{P}$ as a hyperplane section of $Y$, which is $\mathbf{P}=Y\cap\left\{t_{1}=0\right\}$. The fact that $Y$ is not a cone can be proven in the same way as in Lemma 4.2. ∎
@@ 4.5.1 An alternative model for $\mathbf{P}=\mathbf{P}(2,3,10,15)$
$[x:y:z:w]\in\mathbf{P}\mapsto[u^{\prime}:v^{\prime}:s^{\prime}:t^{\prime}]=[x:y^{2}:z:yw].$
@@ 5.3 $\mathbf{P}=\mathbf{P}(1,2,3,6)$
$\varphi^{-1}:[a_{0}:a_{1}:a_{2}:b]\in\mathbf{P}^{\prime}\mapsto[x:y:z:w]=[a_{0}:a_{0}a_{1}:a_{0}^{2}a_{2}:a_{0}^{3}b].$
@@
Consider the open subset $U=\left\{z\neq 0\right\}\subset\mathbf{P}(1,2,3,6)$. By generality of $C$, this opens subset contains the two points $D\cap C$. On the one hand $U$ is isomorphic to the quotient of $\mathbf{C}^{3}_{(\mathtt{x},\mathtt{y},\mathtt{w})}$ under the following action of the group of cubic roots of unity $\mu_{3}$:
@@
On the one hand, we have $h^{0}\mathcal{O}_{\mathbf{P}}(-K_{\mathbf{P}})=27$, while $h^{0}\mathcal{O}_{\mathbf{P}^{\prime}}(5)+3=30$, so even the quintic $f_{5}$ can’t be general, and it must belong to a subspace $V\subset H^{0}(\mathbf{P}^{\prime},\mathcal{O}_{\mathbf{P}^{\prime}}(5))$ with $\dim V=\dim H^{0}(\mathbf{P}^{\prime},\mathcal{O}_{\mathbf{P}^{\prime}}(5))-3=24$. Namely, it follows from the expression of $\varphi^{-1}$ that $f_{5}$ does not involve the monomials $a_{2}^{5},a_{2}^{4}a_{1}$ and $a_{2}^{3}a_{1}^{2}$.

### v3

@@ 1. $\bullet$
    However, these two conditions also make sense for a normal, possibly singular, surface. As already mentioned in the introduction above, the surface $S$ we obtain as a general anticanonical divisor of a Gorenstein weighted projective $3$-space $\mathbf{P}$ is expected to have canonical singularities (to be more precise, ADE singularities). However, it is normal and satisfies the conditions $K_{S}=0$ and $h^{1}(S,\mathcal{O}_{S})=0$. As other authors did before, we refer to such a surface $S$ as a K3 surface with ADE singularities, and we sometimes simply say K3 surface.
@@ 2 The Gorenstein weighted projective spaces of dimension $3$
Remark. If $\mathbf{P}$ is not Gorenstein, then the canonical divisor $K_{\mathbf{P}}$ is not Cartier around at least one coordinate point, i.e., a point given by the vanishing of $3$ out of the $4$ homogeneous coordinates. As an example, consider $\mathbf{P}=\mathbf{P}(1,2,4,5)_{[u:v:s:t]}$, with $\mathcal{O}_{\mathbf{P}}(-K_{\mathbf{P}})=\mathcal{O}_{\mathbf{P}}(12)$. Since $5$ is not a divisor of $12$, then any local section of $\mathcal{O}_{\mathbf{P}}(12)$ around $p_{t}=[0:0:0:1]$ is a quotient of polynomials with numerator a sum of monomials, all of which admit at least one of the coordinates $u,v,s$ as a factor. Hence such a local section vanishes at $p_{t}$.
@@
Moreover, up to isomorphism, we may assume that any three of the four weights $a_{0},a_{1},a_{2},a_{3}$ are coprime. This condition is called well-formedness.
@@ 2.2 The very ampleness of $-K_{\mathbf{P}}$
Let $S$ be a general member of $|-K_{\mathbf{P}}|$ and $\Gamma$ a general member of $|-K_{\mathbf{P}}|_{S}|$. We first prove that $-K_{\mathbf{P}}|_{S}$ is very ample. Thanks to Lemma 2.7, it is enough to prove that $H^{0}(S,\mathcal{O}_{S}(-K_{\mathbf{P}}|_{S}))^{\otimes n}\to H^{0}(S,\mathcal{O}_{S}(-nK_{\mathbf{P}}|_{S}))$ is onto for all $n\geq 1$. It is naturally the case for $n=1$. Assume now that $n>1$ and that it holds up to rank $n-1$. Consider the following commutative diagram
@@
Moreover, we know that K3 surfaces with ADE singularities are projectively normal. Indeed, if $S$ is a singular K3 surface in $\mathbf{P}^{g}$, we may consider a resolution $S^{\prime}\to S\subset\mathbf{P}^{g}$ with $S^{\prime}$ a smooth K3 and apply [SD74, Theorem 6.1.$(ii)$]. It follows that the restriction map $H^{0}(\mathbf{P},\mathcal{O}_{\mathbf{P}}(-nK_{\mathbf{P}}))\to H^{0}(S,\mathcal{O}_{S}(-nK_{\mathbf{P}}|_{S}))$ is onto for all $n\geq 1$. Consider the following commutative diagram with $n>1$, whose right column is exact
@@ 3 Birational models
The first step consists in the introduction of a suitable Veronese map on each $\mathbf{P}$, which is an embedding $v_{n}:\mathbf{P}\hookrightarrow X$, where $X$ is a weighted projective space of dimension $4$, so that the anticanonical model $\mathbf{P}\to\mathbf{P}^{g+1}$ factors as a composition
${\mathbf{P}}$${X}$${\mathbf{P}^{g+1}.}$$\scriptstyle{v_{n}}$
Each time, $X\dashrightarrow\mathbf{P}^{g+1}$ is a rational map that we will specify.
@@ 3.1 The Veronese maps
By a divisibility criterion, we can check fairly easily that $|\mathcal{O}_{X}(\frac{\sigma}{n})|$ is not always basepoint-free. This criterion is purely numerical: $|\mathcal{O}_{X}(\frac{\sigma}{n})|$ is basepoint-free if and only if $\frac{\sigma}{n}$ is divisible by all the weights of $X$ (see this Remark, which is based on the same argument). Namely, it is not the case for #10, #12 and #14, for which the induced map $X\dashrightarrow\mathbf{P}^{g+1}$ is nonregular.
@@ 3.2 The birational models
In all cases but #11, the image of $X$ in $\mathbf{P}^{g+1}$ is a cone with vertex a point which belongs to the image of $\mathbf{P}\to\mathbf{P}^{g+1}$. This follows from the fact that $\frac{\sigma}{n}$ equals the largest weight of $X$; say, $X=\mathbf{P}(d_{0},d_{1},d_{2},d_{3},d_{4})$ with $\frac{\sigma}{n}=d_{4}$. Then the map given by the linear system $|\mathcal{O}_{X}(\frac{\sigma}{n})|$ is the following:
$[x_{0}:x_{1}:x_{2}:x_{3}:x_{4}]\in X\mapsto[\mathtt{f}_{0}:\mathtt{f}_{1}:\cdots:\mathtt{f}_{g}:x_{4}],$
@@
$\varphi:[x:y:z:w]\in\mathbf{P}\mapsto[u_{0}:u_{1}:u_{2}:v]=[x^{5}:xy:z:w]\in\mathbf{P}^{\prime}.$
@@
$u_{0}f_{4}(\mathbf{u},v)+u_{1}^{5}=g_{4}(\mathbf{u},v)=0,$
with $g_{4}$ a general quartic on $\mathbf{P}^{\prime}$. This is summed up in the following commutative diagram
@@ 4 Extensions of $\mathbf{P}$
Let $\Gamma\subset\mathbf{P}^{g-1}$ be a smooth canonical curve of genus $g\geq 11$ such that $\mathrm{Cliff}(\Gamma)\geq 3$. Let $\alpha=\mathrm{cork}(\Phi_{K_{\Gamma}})$ and $Y\subset\mathbf{P}^{g-1+\alpha}$ be a maximal extension of $\Gamma$ with $\dim Y\geq 2$. Then $Y$ is the universal extension of $\Gamma$ if and only if it contains no cone over $\Gamma$ as a linear surface section.
@@
Let us denote by $\mathcal{H}$ the family of $g$-planes $\Lambda\subset\mathbf{P}^{g-1+\alpha}$ such that $\Gamma\subset\Lambda$, and $\mathcal{S}=\mathbf{P}(\mathrm{coker}(\Phi_{K_{\Gamma}}))$ the space which parametrizes the surface extensions of $\Gamma$. Then we have $\mathcal{H}\simeq\mathcal{S}\simeq\mathbf{P}^{\alpha-1}$, and the map $\mathcal{H}\to\mathcal{S}$ which maps $\Lambda$ to $Y\cap\Lambda$ is well-defined if and only if no linear surface section of $Y$ is a cone over $\Gamma$. In that case, it lifts to a linear map $\mathbf{C}^{\alpha}\to\mathbf{C}^{\alpha}$ which has trivial kernel, i.e., it is an isomorphism (see [CD24’, 6.6]).
It follows that $Y$ contains all the surface extensions of $\Gamma$ as unique linear sections, if and only if it contains no cone over $\Gamma$ as linear surface section. In this situation, by Theorem 2.3, it is the universal extension of $\Gamma$. ∎
@@ 4.1 $\mathbf{P}=\mathbf{P}(1,4,5,10)$
The fivefold $Y$ in $\mathbf{P}^{24}$ is a maximal extension of $\Gamma$ by Theorem 2.3 since it has dimension $1+\alpha(\Gamma,K_{\Gamma})$ and it contains $\mathbf{P}=\mathbf{P}(1,4,5,10)$ as a $3$-fold linear section. Indeed, as indicated in Table 3, $\mathbf{P}$ embeds into $\mathbf{P}(1,1,1,2,4)_{[\mathbf{u}:v:s_{0}]}$ as the quintic hypersurface $u_{0}s_{0}=u_{1}^{5}$. It follows that $Y$ is an extension of $\mathbf{P}$, as
@@ 2. $(ii)$
which are all linear sections of $Y$ and contain $\Gamma$ as a hyperplane section. Indeed, the curve $\Gamma$ is cut out on all of them by $\left\{g_{4}(\mathbf{u},v)=0\right\}$. Among them, those that are members of the linear system $\mathcal{L}$ are those parameterized by $\lambda_{1}=\lambda_{2}=0$. The other surface extensions of $\Gamma$ could not be recovered from $\mathbf{P}$ and require the introduction of the birational model $\varphi:\mathbf{P}\dashrightarrow\mathbf{P}^{\prime}$.
@@ 4.2 $\mathbf{P}=\mathbf{P}(1,2,6,9)$
Besides, the embedding of $Y$ in $\mathbf{P}^{30}$ is not a cone, by the same arguments as those mentioned in the proof of Lemma 4.2. The fourfold $Y$ is thus a maximal extension of $\Gamma$. Recall from Table 3 that $\mathbf{P}$ is the hypersurface of $\mathbf{P}(1,1,3,5,9)_{[\mathbf{u}:v:s:t_{0}]}$ given by the equation $u_{0}t_{0}=s^{2}$. This shows that $\mathbf{P}=Y\cap\left\{t_{1}=0\right\}$. ∎
@@ 4.3 $\mathbf{P}=\mathbf{P}(1,3,8,12)$
Hence it is a maximal extension of $\Gamma$. It is also an extension of $\mathbf{P}$; indeed, we know from Table 3 that $\mathbf{P}$ is the hypersurface $u_{0}t_{0}=v^{3}$ is $\mathbf{P}(1,1,3,4,8)_{[\mathbf{u}:v:s:t_{0}]}$. This exhibits $\mathbf{P}$ as a hyperplane section of $Y$, which is $\mathbf{P}=Y\cap\left\{t_{1}=0\right\}$. The fact that the image of $Y$ in $\mathbf{P}^{27}$ is not a cone can be proven in the same way as in Lemma 4.2. ∎
@@ 4.5.1 An alternative model for $\mathbf{P}=\mathbf{P}(2,3,10,15)$
$\psi:[x:y:z:w]\in\mathbf{P}\mapsto[u^{\prime}:v^{\prime}:s^{\prime}:t^{\prime}]=[x:y^{2}:z:yw]\in\mathbf{P}(1,3,5,9).$
@@ 5.3 $\mathbf{P}=\mathbf{P}(1,2,3,6)$
$\varphi^{-1}:[a_{0}:a_{1}:a_{2}:b]\in\mathbf{P}^{\prime}\mapsto[x:y:z:w]=[a_{0}:a_{0}a_{1}:a_{0}^{2}a_{2}:a_{0}^{3}b]\in\mathbf{P}.$
@@
Consider the open subset $U=\left\{z\neq 0\right\}\subset\mathbf{P}(1,2,3,6)$. By generality of $C$, this open subset contains the two points $D\cap C$. On the one hand $U$ is isomorphic to the quotient of $\mathbf{C}^{3}_{(\mathtt{x},\mathtt{y},\mathtt{w})}$ under the following action of the group of cubic roots of unity $\mu_{3}$:
@@
On the one hand, we have $h^{0}(\mathbf{P},\mathcal{O}_{\mathbf{P}}(-K_{\mathbf{P}}))=27$, while $h^{0}(\mathbf{P}^{\prime},\mathcal{O}_{\mathbf{P}^{\prime}}(5))+3=30$, so even the quintic $f_{5}$ can’t be general, and it must belong to a subspace $V\subset H^{0}(\mathbf{P}^{\prime},\mathcal{O}_{\mathbf{P}^{\prime}}(5))$ with $\dim V=\dim H^{0}(\mathbf{P}^{\prime},\mathcal{O}_{\mathbf{P}^{\prime}}(5))-3=24$. Namely, it follows from the expression of $\varphi^{-1}$ that $f_{5}$ does not involve the monomials $a_{2}^{5},a_{2}^{4}a_{1}$ and $a_{2}^{3}a_{1}^{2}$.